OpenStudy (anonymous):

AP CALCULUS BC Find the Maclaurin polynomial P4 for f(x) = x^2 e^x. I have the answer, but I would like to see the method of getting it.

OpenStudy (anonymous):

\[f(x) = f(0) + x f'(0) + \frac{x^2 f''(0)}{2!} + \frac{x^3 f'''(0)}{3!} + \frac{x^4 f''''(0)}{4!}+...\]

OpenStudy (anonymous):

this is the maclaurin series for any function f(x) here \[f'(0) = \frac{d f(x)}{dx} |_{x= 0}\]

OpenStudy (anonymous):

P4 is the polynomial till the fourth derivative as shown in the sample equation.

OpenStudy (anonymous):

differentiate the given function four times ( to get the four derivatives needed ) and plug x = 0 into the obtained derivatives and then substitute those derivatives in the place of f'(0), f''(0), f'''(0) and f''''(0)

OpenStudy (anonymous):

any doubts?

OpenStudy (anonymous):

Ok, I'm going to do this now on paper and see if I get it.

OpenStudy (anonymous):

Ohhhh, ok. I got it! Parts of the derivatives start cancelling out. I was wondering what happened to the factorials in the denominator lol. Thanks for clearing it up :)

OpenStudy (anonymous):

Its been a pleasure :)